OpenStudy (anonymous):

Write the equation of a hyperbola with vertices at (1,1) and (9, 1) and foci at (0, 1) and (10, 1).

OpenStudy (anonymous):

ok here we go what is the center? it is half way between \((1,1)\) and \((9,1)\)

OpenStudy (anonymous):

The length of the transverse axis is the distance between the vertices: 2a = 8, a = 4. Since the vertices are on a horizontal line, y = 1, the transverse axis is horizontal, and the hyperbola is of the form: (x-g)²/4² - (y-h)²/b² = 1 The center is halfway between the vertices: (g, h) = (10/2, 2/2) = (5, 1) (x-5)²/4² - (y-1)²/b² = 1 The distance between the foci = 2c = 10 => c = 5 c² = a² + b² 25 = 16 + b² b = 3 The equation of the hyperbola, in "conic" form is: (x-5)²/4² - (y-1)/3² = 1 The equation of the hyperbola, in standard form is: 9x² - 16y² - 90x + 32y + 65 = 0

OpenStudy (anonymous):

you got this, or you want to do it step by step? it is not hard

OpenStudy (anonymous):

I think I have it (: thanks

OpenStudy (anonymous):

yw